A pendulum consists of a small spherical mass suspended from a rigid
support by a length of thread of negligible mass.
The mass is swinging from side to side with an amplitude of 0.05 m and
a period of 1.40 s.
The motion of the mass can be considered to be simple harmonic.
Calculate the maximum acceleration of the mass.
1
Expert's answer
2020-12-14T07:16:40-0500
a=ω2A=4π2T2Aa=4π21.420.05=1.01ms2a=\omega^2A=\frac{4\pi^2}{T^2}A\\ a=\frac{4\pi^2}{1.4^2}0.05=1.01\frac{m}{s^2}